# Are LLMs Good Literature Review Writers? Evaluating the Literature Review Writing Ability of Large Language Models

### v2

@@ Abstract
The literature review is a crucial form of academic writing that involves complex processes of literature collection, organization, and summarization. The emergence of large language models (LLMs) has introduced promising tools to automate these processes. However, their actual capabilities in writing comprehensive literature reviews remain underexplored, such as whether they can generate accurate and reliable references. To address this gap, we propose a framework to assess the literature review writing ability of LLMs automatically. We evaluate the performance of LLMs across three tasks: generating references, writing abstracts, and writing literature reviews. We employ external tools for a multidimensional evaluation, which includes assessing hallucination rates in references, semantic coverage, and factual consistency with human-written context. By analyzing the experimental results, we find that, despite advancements, even the most sophisticated models still cannot avoid generating hallucinated references. Additionally, different models exhibit varying performance in literature review writing across different disciplines.
@@ 1 Introduction
The literature review is a form of academic writing designed to summarize, analyze, and evaluate existing research findings, theoretical perspectives, and methodologies. However, writing a literature review often requires extensive reading and summarizing of relevant literature. Literature reviews in popular fields may require tens or even hundreds of references, which is a complex and labor-intensive task. In recent years, deep learning models have been developed to assist in the automated writing of literature reviews Kontonatsios et al. (2020); Aliyu et al. (2018). The advent of large language models (LLMs) has brought promising advances in automating certain stages of literature review writing. With their powerful generative capabilities, enhanced by training on vast amounts of academic text, LLMs have emerged as a formidable tool for writing literature reviews Wang et al. (2024).
However, how well LLMs perform in writing literature reviews remains unclear. To the best of our knowledge, no comprehensive studies have yet evaluated the literature review writing capabilities of LLMs. Literature review writing requires, on one hand, deep familiarity with the studies in a specific field, and on the other hand, the ability to summarize the core contributions of that literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance LLMs’ domain-specific knowledge, most researchers use LLMs without RAG such as ChatGPT for literature review writing. Therefore, it is necessary to evaluate the literature review writing capabilities of these native LLMs.
Therefore, in this paper, we propose a framework for automatically assessing literature review writing ability in LLMs. The evaluation of LLMs’ ability to write literature reviews remains an open question. In this study, we use human-written literature reviews as the gold standard and employ external tools for a comprehensive evaluation. Specifically, we first collect human-written literature reviews to create the dataset. Then, we ask LLMs to complete three tasks: generating references, writing an abstract, and writing a literature review based on a given topic. The generated results are then evaluated from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context.
@@
  • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement. This framework encompasses multiple stages, including dataset construction, LLMs gnenrated data collection, and evaluation.
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks: generating references, writing an abstract, and writing a literature review based on a given topic.
  • Then, we evaluate the generated results of the LLMs from multiple aspects, including the hallucination rate in generated references, factual consistency with human-written results, and the comparison of semantic coverage.
  • Finally, we assess four LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain an unavoidable issue for current LLMs. Additionally, the performance of LLMs in writing literature reviews varies across different disciplines.
@@ 2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024) proposed a literature survey writing process for LLMs, called AutoSurvey. By RAG retrieving the latest relevant paper incorporates real-time knowledge. In evaluation, Muti-LLM-as-judge is used to evaluate citation quality and content quality.  Wang_Hu_Wang_Yan_Sheng_He_2024 compared the use of LLMs across different academic tasks and found that researchers are more likely to use LLMs for literature review (LR) tasks than for paper understanding (PU) tasks.
Additionally, more attention has been paid to the hallucination evaluation of reference citations. According to whether the output contradicts the prompt or not,  OpenAI et al. (2023) categorized hallucination phenomenon as open domain and closed-domain hallucinations. Closed-domain hallucination assessments require only comparisons with input contexts, whereas open-domain hallucination assessments require more extensive research Bubeck et al. (2023), such as human evaluation Ji et al. (2023).
Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generated 11 systematic reviews under the shoulder rotator cuff pathology topic, then to evaluate the hallucination rates of reference. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. For example,  Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rate. They first asked LLMs to generate 5 reference titles, then use BING Search API to label the reference automatically as true or not. The LLMs were then examined to see if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?).
@@
Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on the given 50 topics. ChatGPT generated 178 references in these proposals and then they employed five researchers to manually check these references and DOIs on Scopus, Google and PubMed search engines and found that among the references 109 references had valid DOIs. score  Aljamaan et al. (2024) given 5 medical topics, then used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generated 10 relevant references for each topic, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication data, link to the paper, and relevance of the reference to the keyword prompt. Then, they proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. The existing studies on evaluating citation hallucinations in literature reviews are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
@@ 3 Methodology
In this section, we present the framework for evaluating LLMs’ literature survey writing ability. The framework as shown in Figure 1, consists of three main stages: first, dataset construction; second, task design for evaluation; and third, assessment of the generated results.
@@ 3.1 Dataset Construction
As illustrated in Figure 1, we first collect literature survey data from the Annual Reviews website ( https://www.annualreviews.org/). Annual Reviews is an independent, nonprofit scholarly publishing company, which publishes 51 journals of review articles. The articles on the website are peer-reviewed, so they are of high quality and can be used as experimental materials. We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references. Then clean them as the experimental dataset. Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex. In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent.
Then, the dataset $D$ is the article set from $K$ journals, $D=\{j_{0},...,j_{k},...,j_{K}\}$, the number of journals is 51, so $K=50$. For each journal $j_{k}$, there is a set of articles $j_{k}=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles in $j_{k}$, and each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{l},...,r_{L}\}$.
@@ 3.2 Task Design
We design three tasks as follows.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract according to the research topic provided in the title and the keywords. The length of the generated abstract is required to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a literature review according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@ 3.3 Evaluation Metrics
Figure 1: Illustration of the proposed framework.
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate accurate references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM.
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task1 and Task3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $C_{t}$ and $C_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $C_{t}$ and $C_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $c_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $c_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. therefore, when the other three elements in $r_{n}^{g}$ are correct, we also consider it as a reliable article.
@@
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, for each journal $j_{k}$, we can obtain the accuracy of generated references $Acc_{ref.}(j_{k})$ based on Eq 3. Subsequently, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(j_{k})$ across journals, as shown in Eq 4. Finally, we defined the reference accuracy metric $Acc$ as Eq 4. A higher accuracy metric indicates a lower hallucination rate.
@@
$Acc_{ref.}(j_{k})=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
$Acc.=\frac{1}{K}\sum_{k=0}^{K}{Acc_{ref.}(j_{k})}$ (4)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of each journal and each LLM, formulated as follows.
$S_{t}=1/(NMK)\sum_{K}\sum_{M}\sum_{N}s_{p_{i}}$ (5)
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from Semantic Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (6)
where $s$ is the number of correct generated reference titles for $p_{i}$. We search the generated title from an external academic search engine to estimate whether the title can be matched.
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{i}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
$Entail_{i}=\theta_{\text{NLI}}(a_{i},a_{i}^{g})=\begin{cases}1&\text{if}a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (7)
@@
$Entail=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}{Entail_{i}}$ (8)
@@
KPR, first proposed by  Qi_Xu_Guo_Wang_Zhang_Xu_2024, is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents. Since the human-written literature review is lengthy and difficult to compare directly, we adapt the KPR metric by first using GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculating the coverage of these key points by the model-generated literature review as Eq. 9.
$KPR=\frac{1}{KM}\sum_{k=0}^{K}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (9)
where $c_{i}^{g}$ denotes the literature review generated by LLMs. Finally, we also concatenate key points and compute the ROUGE metric between the key point and $c_{i}^{g}$.

### v3

# Large Language Models for Automated Literature Review: An Evaluation of Reference Generation, Abstract Writing, and Review Composition

@@ Abstract
Large language models (LLMs) have emerged as a potential solution to automate the complex processes involved in writing literature reviews, such as literature collection, organization, and summarization. However, it is yet unclear how good LLMs are at automating comprehensive and reliable literature reviews. This study introduces a framework to automatically evaluate the performance of LLMs in three key tasks of literature writing: reference generation, literature summary, and literature review composition. We introduce multidimensional evaluation metrics that assess the hallucination rates in generated references and measure the semantic coverage and factual consistency of the literature summaries and compositions against human-written counterparts. The experimental results reveal that even the most advanced models still generate hallucinated references, despite recent progress. Moreover, we observe that the performance of different models varies across disciplines when it comes to writing literature reviews. These findings highlight the need for further research and development to improve the reliability of LLMs in automating academic literature reviews.
@@ 1 Introduction
The literature review is a critical component of academic writing that aims to synthesize, critique, and assess the current state of knowledge in a particular field. It involves a comprehensive examination of published research articles, theoretical frameworks, and research methodologies related to a specific topic. Conducting a thorough literature review often necessitates extensive reading and summarizing of pertinent literature, which can be a complex and time-consuming process, especially in well-established fields where the number of relevant references can range from dozens to hundreds. To alleviate this burden, researchers have recently turned to advanced deep learning models as a potential tool to aid in the automated generation of literature reviews. Kontonatsios et al. (2020); Aliyu et al. (2018). The emergence of large language models (LLMs) has introduced a promising avenue for automating various aspects of the literature review writing process. LLMs, trained on extensive collections of academic texts, have developed impressive generative capabilities that make them a potent tool for assisting researchers in crafting literature reviews. By leveraging the knowledge acquired from vast amounts of scholarly literature, these models can help researchers quickly identify relevant sources, summarize key findings, and generate coherent syntheses of the existing body of work Wang et al. (2024b); Agarwal et al. (2024).
The process of writing a literature review requires a deep understanding of the studies within a particular field, as well as the ability to summarize the core contributions of the relevant literature. While techniques such as Retrieval-Augmented Generation (RAG) can enhance the domain-specific knowledge of LLMs, most researchers rely on vanilla LLMs, such as ChatGPT, for literature review writing without the use of RAG Wang et al. (2024a). Consequently, it is crucial to evaluate the performance of these native LLMs in the context of literature review writing to determine their effectiveness and limitations.
Therefore, in this paper, we propose a framework for automatically assessing the literature review writing ability of LLMs, using human-written literature reviews as the gold standard and designing metrics for a comprehensive evaluation. We first collect a dataset of human-written literature reviews to serve as a benchmark for evaluating the performance of LLMs. We then ask LLMs to complete three tasks based on the collected dataset: generating references, writing an abstract, and writing a complete literature review based on a given topic. Finally, we evaluate the generated results from several dimensions, including the presence of hallucinations in the references, as well as the semantic coverage and factual consistency of the generated literature review compared to the human-written context. By assessing the performance of LLMs across these tasks and evaluating their output using our proposed metrics, we aim to provide a comprehensive understanding of their capabilities and limitations in writing literature reviews.
@@
  • First, we propose a framework for automatically evaluating the literature review writing ability of LLMs, without requiring any human involvement. This framework encompasses multiple stages, including the compilation of a literature review dataset construction, the collection of LLM-generated output, and the evaluation of LLM performance.
  • Second, we collect 1,105 literature reviews from 51 journals across 5 disciplines as the ground truth. We then design three tasks for accessing LLMs in literature writing: reference generation, literature summary, and literature composition on a given topic.
  • Then, we evaluate the generated results of LLMs from multiple perspectives, including the hallucination rate in generated references, factual consistency, and semantic coverage compared to human-written content.
  • Finally, we assess five LLMs using the proposed framework. By analyzing the experimental results, we find that hallucinated references remain a prevalent issue for current LLMs. Furthermore, the performance of LLMs in writing literature reviews varies across different disciplines.
@@ 2 Related Work
Recent studies have increasingly focused on using LLMs for writing literature reviews. For example,  Wang et al. (2024b) proposed a literature survey writing process for LLMs, called AutoSurvey, which incorporates real-time knowledge by retrieving the latest relevant papers using a retrieval-augmented generation (RAG) approach. Similarly,  Agarwal et al. (2024) investigated the zero-shot capabilities of recent LLMs in generating literature reviews from abstracts. They employed a two-step approach: first retrieving relevant papers via an external knowledge base, then outlining and executing a structured plan to generate the review. The authors focused on the LLMs’ ability to effectively summarize the literature.
@@
Additionally, for evaluating the literature review writing capabilities of LLMs, more attention has been paid to assessing hallucinations in reference citations. For example, Chelli_Descamps_Lavoué_Trojani_Azar_Deckert_Raynier_Clowez_Boileau_Ruetsch-Chelli_2024 used LLMs (ChatGPT, GPT-4, and Bard LLM) to generate 11 systematic reviews on shoulder rotator cuff pathology and evaluated the hallucination rates of the references. They found that Bard had significantly higher rates of hallucinations than ChatGPT and GPT-4. Similarly, Agrawal et al. (2024) chose 200 computer science topics to evaluate LLMs’ hallucination rates. They first asked LLMs to generate 5 reference titles, then used the BING Search API to automatically label the references as true or not. The LLMs were then examined to determine if they could actively discover the hallucinated literature through direct queries (asking directly if the paper exists) and indirect queries (who wrote the paper?). Athaluri et al. (2023) assessed the hallucinations of LLMs in research proposal writing. They used ChatGPT to generate 50 research proposals based on 50 given topics. ChatGPT generated 178 references in these proposals, and then they employed five researchers to manually check these references and DOIs on Scopus, Google, and PubMed search engines. They found that among the references, 109 had valid DOIs. Moreover, Aljamaan et al. (2024) provided 5 medical topics and used ChatGPT 3.5, Bard, Perplexity, Bing, Elicit, and SciSpace to generate 10 relevant references for each topic, with each reference including 7 items: title, journal name, author, DOI, publication date, link to the paper, and relevance of the reference to the keyword prompt. They then proposed the Reference Hallucination Score (RHS) for citation hallucination assessment of LLMs, assigning different hallucination scores to each item, e.g., a score of 2 corresponding to the title and 1 corresponding to the publication date. While these studies are very useful attempts at evaluating citation hallucinations in literature reviews, they are limited to specific academic topics and do not provide a comprehensive assessment of LLM performance. Moreover, these evaluations often involve manual human intervention.
@@ 3 Methodology
In this section, we propose a framework for evaluating LLMs’ literature survey writing ability. The framework, as shown in Figure 1, consists of three main stages: dataset construction and task design for evaluation, collection LLM-generated output, and assessment of the generated output.
@@ 3.1 Dataset Construction
Assessing the ability to write literature reviews is a challenging task, as evaluating the quality of content is inherently complex. In this paper, we use human-written reviews as the gold standard, which simplifies the evaluation process to some extent. As illustrated in Figure 1, we first collect publicly available information of literature reviews (i.e., the title, authors, abstract, and keywords) from the Annual Reviews website (https://www.annualreviews.org/). Annual Reviews, an independent nonprofit publisher, produces 51 review journals spanning various scientific disciplines. Invited experts write comprehensive, authoritative reviews that synthesize and summarize the most significant primary research literature in their field, providing a valuable resource for researchers to stay current with the latest developments. We crawl all articles published in 2023, including their title, keywords, abstracts, contents, and references, and then clean them to create the experimental dataset.
Then, the dataset $D$ is the article set from 51 journals, $D=\{p_{0},...,p_{i},...,p_{M}\}$, where $M$ represents the number of articles. Each article $p_{i}=\{t_{i},w_{i},a_{i},c_{i},R_{i}\}$, where $t_{i}$, $w_{i}$, $a_{i}$, $c_{i}$, $R_{i}$ represent the title, keywords, abstract, context, and reference set $R_{i}=\{r1,...,r_{k},...,r_{K}\}$, and $K$ represents the length of the reference set.
@@ 3.2 Task Design
Since literature review writing primarily involves the collection and synthesis of relevant research, we design three tasks as follows to evaluate LLMs’ capabilities in different aspects of literature review writing.
  • Task 1: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to find the $N$ most relevant studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ to the research topic. Each citation study must include 7 metadata elements: title, authors, journal, year, volumes, first page, and last page. In this task, we evaluate whether LLMs can recommend reliable references based on the given topic.
  • Task 2: Given the article title $t_{i}$ and keywords $w_{i}$, ask LLMs to write an abstract $a_{i}^{g}$according to the research topic provided in the title and the keywords. The length of the generated abstract is to be the same as the length of the original abstract of the article. In this task, we evaluate whether LLMs can write an abstract based on the given topic.
  • Task 3: Given the article $t_{i}$ and keywords $w_{i}$, and abstract $a_{i}$, ask LLMs to write a short literature review $c_{i}^{g}$ according to the research topic provided in the title, keywords, and abstract. The literature survey should be about 1000 words long. LLMs also need to back up claims by citing previous studies $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ (with a total of $N$ citations in the literature survey). In this task, we evaluate whether LLMs can write a high-quality literature survey and cite truth studies.
@@ 3.3 Evaluation Metrics
Figure 1: Illustration of the experiment process.
Reference hallucination evaluation metrics. Given that LLMs are trained on vast corpora, including academic sources, we aim to evaluate whether they can generate true references. In this section, we introduce the calculation process of the reference accuracy $Acc.$ and title search rate $S_{t}$ for each LLM. A higher accuracy metric indicates a lower hallucination rate.
For each article $p_{i}\in D$, each LLM generates $N$ references $R_{i}^{g}=\{r_{1}^{g},...,r_{n}^{g},...,r_{N}^{g}\}$ in Task 1 and Task 3, each $r_{n}^{g}=\{T,A,J,Y,V,FP,LP\}$ and includes 7 elements corresponding to the generated article title, authors, journal, year, volume, first page, and last page. Each element corresponding to a state label represents whether it is accurate or not $\{e_{d}\mid d=0,1,\dots,7\}$, $e_{d}$ = 1 or 0.
Next, we describe how to obtain $\{e_{d}\mid d=0,1,\dots,7\}$. First, we use the generated titles $T$ and the first author in $A$ as the queries and search them separately from external academic search engines. Then, we obtain two candidate sets $Z_{t}$ and $Z_{a}$ respectively. Subsequently, we compare the generated $r_{n}^{g}$ with the article from candidate sets $Z_{t}$ and $Z_{a}$. For example, if the title of the article candidate is the same as that of $r_{n}^{g}$, then $e_{0}=1$. Finally, we find the best candidate article based on the sum of $\{e_{d}\mid d=0,1,\dots,6\}$, and the one with the largest sum is the best candidate article $z_{n}^{g}$ of $r_{n}^{g}$.
Then, we compare the algin degree between the generated $r_{n}^{g}$ and the best candidate article $z_{n}^{g}$ to determine whether $r_{n}^{g}$ corresponds to a true article (As shown in Eq 1). On the one hand, when its title $T$ is correct and one other element is correct, we consider $r_{n}^{g}$ corresponds to a true article. Here its title $T$ is correct or not, we give LLM a certain margin of error, as long as the title has a match rate of 80%, it will be judged as accurate, so we need another piece of information that is correct to ensure that the article is reliable. On the other hand, when its title $T$ is incorrect, certain other information is correct which can also help us to locate the article. Therefore, when the other three elements in $r_{n}^{g}$ are matching, we also consider it as a reliable article.
@@
For each paper $p_{i}$ in the dataset, we formulate the accuracy of the references generated by the LLM as Eq 2. Then, we obtain the accuracy of generated references for each LLM $Acc.$ by averaging $Acc_{ref.}(p_{i})$ across articles, as shown in Eq 3.
@@
$Acc.=\frac{1}{M}\sum_{i=0}^{M}Acc_{ref.}(p_{i})$ (3)
On the other hand, intuitively, the title is the key element that reflects the generated reference’s faithfulness. In  Agrawal et al. (2024)’s work, they determined by ground-truth labels assigned using the Bing search API. Encouraged by this work, we also calculate the title search scores of generated references for each LLM, formulated as follows.
$S_{t}=\frac{1}{MN}\sum_{M}\sum_{N}s_{p_{i}}$ (4)
$s_{p_{i}}=\sum_{r_{n}^{g}\in R_{i}^{g}}\begin{cases}1&\text{if the $T\in r_{n}^{g}$ has return value }\\ &\text{from external Scholar API,}\\ 0&\text{otherwise}\end{cases}$ (5)
where $s_{p_{i}}$ is the number of the generated title that can be matched in the external scholar base. We search the generated title from an external academic search engine to estimate whether the title can be matched.
Context evaluation metrics. In this paper, we use the human-written article as the gold truth and then evaluate LLM-generated context from factual consistency and semantic coverage aspects. The resemblance of natural language inference (NLI) to factual consistency evaluation has led to utilizing NLI models for measuring factual consistency Gao et al. (2023). Encouraged by previous works, we also use the NLI method to evaluate the factual consistency between LLMs generated and human-written text. For example, we calculate the NLI score $Entail_{p_{i}}$ between the original article abstract $a_{i}$ and the LLM-generated abstract $a_{i}^{g}$ as follows.
$Entail_{p_{i}}=\theta_{\text{NLI}}(a_{i}^{g},a_{i})=\begin{cases}1&\text{if }a_{i}^{g}\text{ entails }a_{i}\text{,}\\ 0&\text{otherwise}\end{cases}$ (6)
@@
$Entail=\frac{1}{M}\sum_{i=0}^{M}{Entail_{p_{i}}}$ (7)
@@
KPR, first proposed by  Qi et al. (2024), is a metric designed to evaluate the effectiveness of LLMs in utilizing RAG for long documents. Since human-written literature reviews are lengthy and difficult to compare directly, we adopt the KPR method to measure the extent to which LLM-generated content covers the key points in human-written literature reviews. Specifically, we first use GPT-4 to extract $q$ key points $X_{i}=[x_{i_{1}},x_{i_{2}},...,x_{i_{q}}]$ from the human-written literature review $c_{i}$, and then calculate the coverage of these key points by the model-generated literature review as Eq. 8.
$KPR=\frac{1}{M}\sum_{i=0}^{M}\frac{\sum_{x\in X_{i}}\theta_{\text{NLI}}(x,c_{i}^{g})}{|X_{i}|}$ (8)
where $c_{i}^{g}$ denotes the literature review generated by LLMs.
Finally, we also concatenate key points and compute the ROUGE metric between the key points and $c_{i}^{g}$.